OpenStudy (anonymous):

what is the slope and y-intercept for y+5x=-7

OpenStudy (anonymous):

put the equation in to y = mx+b form and m would be the slope b would be the y - intercept

OpenStudy (issy14):

y+5x=-7 y=-5x-7 m (slope) = -5 y-intercept: (0,-7)